oregonduck (oregonduck):

2x^+4x+1+4 FACTORING BY GROUPING

pooja195 (pooja195):

\[\huge~\rm~2x^2+4x+1+4?\]

pooja195 (pooja195):

is it like that?

oregonduck (oregonduck):

YES

pooja195 (pooja195):

Combine Like Terms: =2x2+4x+1+4 =(2x2)+(4x)+(1+4) =2x2+4x+5

oregonduck (oregonduck):

no factor like grouping is what i am trying to do

oregonduck (oregonduck):

factoring by grouping

pooja195 (pooja195):

Then we do this 

oregonduck (oregonduck):

?

oregonduck (oregonduck):

just tell be the steps to find the answer

pooja195 (pooja195):

have you ever seen this ? \(\huge\color{purple}{ax^2+bx+c}\)

oregonduck (oregonduck):

yes

pooja195 (pooja195):

That is known as standered form when factoring by grouping we multiply a and c

pooja195 (pooja195):

have you ever seen this ? \(\huge\color{red}{2x^2+4x+5}\)

oregonduck (oregonduck):

yes

pooja195 (pooja195):

lol my typos >_< ok multiply... \[\huge~\rm~5 \times 2=?\]

OpenStudy (anonymous):

10

OpenStudy (anonymous):

XD

oregonduck (oregonduck):

10

OpenStudy (hybrik):

@pooja195 He is asking by grouping

OpenStudy (hybrik):

@OregonDuck You asked this question before

OpenStudy (anonymous):

pooja the english chat nowwwwww

oregonduck (oregonduck):

@mathmate said that that answer is wrong so i am trying to get it checked

OpenStudy (hybrik):

@OregonDuck It was wrong before because you said a different terms

pooja195 (pooja195):

Actually this isnt even facotralbe

OpenStudy (hybrik):

This question can be 2x(x+2) and 5

pooja195 (pooja195):

It would be prime :/ @hybrik i know that he was asking to do it by grouping this is that way i was taught.

OpenStudy (hybrik):

Im getting (2x+5)(x+2)

OpenStudy (hybrik):

@pooja195 I think its prime too, let me check

pooja195 (pooja195):

its prime you cant factor this

OpenStudy (hybrik):

As I checked, you can go to 2x2+4x+5, but, its prime,

pooja195 (pooja195):

Right so @OregonDuck its prime which means not factor-able

oregonduck (oregonduck):

ok